Carbon12 (C12) is stable and is the most abundant isotope of carbon. An ancient scrap of paper has 0.7 times the Caron14 that a new piece of paper has. How many years old is the scrap of paper? (half-life of C14 is 5730 years)


1
Expert's answer
2021-11-26T04:05:02-0500

Carbon dating yt =ln (C14)0 / (C14)t


where (C14)0 is equivalent to amount of carbon 14 in new paper = 1

(C14)t is amount of carbon 14 in an ancient scrpt of paper = 0.7


t = time of decay

y = decay constant

= Ln2/t 1/2

Half life of carbon 14 = 5730

= Ln2/5730 years

Combining all values

(Ln2 x t)/5730 = Ln 1/0.7

t = 5730/Ln2 x Ln1/0.7

=4092.85 years


Answer= 4093 years